# Automatic and Structure-Aware Sparsification of Hybrid Neural ODEs

### v1

@@ 1 Introduction
In the field of dynamical system modeling, a significant category of hybrid approaches builds on neural ordinary differential equations (neural ODEs)  (Haber and Ruthotto, 2017; Chen et al., 2018; Kidger, 2021), which arose from the insight that deep residual neural networks can be formulated as continuous-time dynamical systems (Rico-Martinez et al., 1992; Weinan, 2017). Neural ODEs are well-suited for modeling dynamical systems as they offer continuous-time representations with latent dynamics that integrate seamlessly into modern machine learning pipelines with automatic differentiation, making them both scalable and flexible. More recently, researchers have adapted neural ODEs to incorporate domain-knowledge-informed relational inductive bias, often derived from mechanistic models or causal graphs. This hybrid style—sometimes termed Graph Neural ODE (Poli et al., 2019), Neural Causal Model (Xia et al., 2021), Neural State Space Modeling (Hussain et al., 2021) or Mechanistic Neural ODE (MNODE) (Zou et al., 2024)—ensures mechanistic plausibility and interpretability while taking advantage of the flexibility of neural networks, thereby improving model performance and robustness in data-scarce settings.
@@
Traditional model reduction approaches in biochemistry—such as timescale separation (Michaelis and Menten, 1913; Johnson and Goody, 2011) and quasi-steady-state approximations (Schauer and Heinrich, 1983; Bothe and Pierre, 2010)—often require deep domain expertise or extensive trial-and-error. On the other end of the spectrum, data-driven graph-based model reduction offers a pathway to solve this problem. In many healthcare application domains, mechanistic ODEs can be represented as reaction networks or directed graphs, where nodes denote system states and edges denote interactions (Hodgkin and Huxley, 1952; Holz and Fahr, 2001; Smith et al., 2004; Canini and Perelson, 2014; Man et al., 2014). In recent years, many solutions have emerged from the graph neural network (GNN) community for general graph pruning, using approaches such as topology-based node/edge selection (Spielman and Srivastava, 2008; Liu et al., 2023), learning-based sub-graph sampling (Wang et al., 2019; Zeng et al., 2019; Zheng et al., 2020), or optimization-based graph sparsification (Li et al., 2020; Jiang et al., 2021, 2023). However, these reduction methods are typically data-driven and agnostic of any domain knowledge, and thus do not necessarily preserve key mechanistic structures or constraints. Furthermore, non-gradient-based reduction methods (e.g., greedy search) can be prohibitively costly in computation for large, high-dimensional ODE systems. As a result, a gap remains for computationally efficient solutions that reduce model complexity while preserving the mechanistic integrity and improving predictive performance for hybrid neural ODEs.
@@ 2.1 Preliminary
where $S_{\text{pa}(i)}\subseteq S,\,\,X_{\text{pa}(i)}\subseteq X$ are subsets of state and input variables, respectively, on which the derivative of $s_{i}$ with respect to $t$ depends, i.e., they are “parents" of $s_{i}$.
@@
Observable states are state variables in the system that can be directly measured through experiments or sensors. These are the quantities that can be collected and tracked over time. On the other hand, latent states are state variables that are not directly accessible but still believed to play a role in system dynamics.
@@
2.2 Time series forecasting with dynamic covariates
@@
we would like to predict $S_{\text{obs}}^{\text{F}}=\{S_{\text{obs}}(t_{k})\}_{k=1}^{q}\in\mathbb{R}^{q\times|S_{\text{obs}}|}$, the value of the observable states at future time stamps.
@@
2.3 Mechanistic neural ODE (MNODE)
@@
$\frac{ds_{i}(t)}{dt}=\text{NN}_{i}(S_{\text{pa}(i)}(t),X_{\text{pa}(i)}(t),t).$ (1)
@@
2.4 Hybrid graph sparsification (HGS) for MNODE
Step 1: merging maximally strongly connected components Given a directed graph representation of a mechanistic ODE system $G=(E,V)$ (since the dependency on the mechanistic ODE system is clear from context, we will omit the $M$ subscript to simplify notations), we first collapse all maximal strongly connected components (MSCCs) in $G$ into super-nodes to make it an RDAG (note that we allow self loops). This is implemented by first partitioning $V$ into disjoint subsets of MSCCs $C_{i}$:
@@ 2.4 Hybrid graph sparsification (HGS) for MNODE
Rationale of step 1 Transforming the original graph into an RDAG reveals essential causal structure and provides a topological ordering that improves training stability. Replacing each MSCC with a self-loop allows neural networks to flexibly model complex intra-component dynamics—a key principle of hybrid modeling.
@@
$D_{x,s}=\{v\in V^{a}\mid v\neq x,v\neq s,\,\,\text{$s$ no longer reachable from $x$ in $G^{a}$ after removing $v$}\},$
@@
Rationale of step 2 Mechanistic models often group biological processes into compartments or pathways whose internal dynamics can be approximated with fewer latent states. For example, quasi-steady-state approximations in chemical kinetics eliminate fast variables by assuming equilibrium. In data-scarce settings, such simplifications often reduce variance more than they introduce bias. To support this, we augment the RDAG with the partial transitive closure of essential input–output paths, enabling the model to choose simpler routes while preserving reachability from Step 1. Crucially, using a partial (rather than full) closure prevents introducing direct input–output edges unsupported by the mechanistic model, helping preserve the integrity of latent dynamics.
@@ 3 Related work
Sparsity, LASSO, and Generalization LASSO-style penalties have long been used to induce sparsity. Applications range from residual networks (Lemhadri et al., 2021), varying-coefficient models (Thompson et al., 2023), and CNNs via group LASSO (Liu et al., 2015; Wen et al., 2016), to feature selection in MLPs (Zhao et al., 2015; Sun et al., 2016; Wang et al., 2017) and local linear sparsity (Ross et al., 2017). However, these methods often yield sparsity patterns that are hard to interpret. Our approach extends LASSO to MNODE while grounding the learned sparsity in mechanistic graph structures, providing interpretability and domain-aligned insights.
Prior work has shown that reduced neural networks can generalize as well as, or even better than, their unreduced counterparts (Gale et al., 2019; Bartoldson et al., 2020; Hoefler et al., 2021). For graph-based model reduction, You et al. (2020) found that mildly sparse relational graphs often outperform fully connected ones. Our results support this: MNODEs built from lightly pruned graphs achieve better performance than those using dense graphs in low-data regimes.
@@
Graph Sparsification via Optimization Our method is related to optimization-based graph sparsification approaches for GNNs. For example, Li et al. (2020) constrained the $L_{0}$ norm of the adjacency matrix, Jiang et al. (2021) used elastic net penalties, and Jiang et al. (2023) applied exclusive group LASSO to encourage neighborhood sparsity. Unlike these methods, we do not sparsify the adjacency matrix directly. Instead, we sparsify edge weights in message passing, reducing influence from less informative neighbors.
Crucially, GNN sparsification methods are typically data-driven and ignore mechanistic structure. In contrast, our method is structure-aware: it begins with a domain-informed pruning step that restricts the search to physically plausible graphs. This makes our approach more suitable for hybrid models like the Graph Network Simulator (Sanchez-Gonzalez et al., 2020) that require mechanistic consistency.
Finally, while Zou et al. (2024) used a greedy, stepwise reduction scheme, our method is more computationally efficient and yields more stable predictive performance—analogous to the gains of LASSO over stepwise selection in linear models (Hastie et al., 2020).
@@ 4.1 Experiments on synthetic data
Starting graph Assuming the true data generating process is unknown, we consider two settings in which the starting “mechanistic" graph contains redundant structures: (1) a refined graph whose redundant part contains 3 input nodes, 1 latent node, and 1 latent cycle; (2) a comprehensive graph whose redundant part contains 6 input nodes, 3 latent nodes and 3 latent cycles (See Appendix A3.1 for illustrations). All redundant input variables are generated from independent $\mathcal{N}(0,0.5)$.
@@
Evaluation and metrics All evaluations are performed on the test set (see Appendix A4). We report RMSE with 1-sigma standard error (SE) for predictive performance, Peak (worst-case) RMSE for robustness and Effective Number of Parameters (ENP, average number of parameters whose magnitude is $\leq 10^{-3}$ under CV-selected hyperparameter setting) for model reduction. Additional metrics including MAPE, Peak MAPE and correlation are reported in A6.1.
@@ 4.2 Experiments on real-world data: blood glucose forecasting for T1D patients
Data Our data are from the T1D Exercise Initiative (T1DEXI)  (Riddell et al., 2023), which is available to public at https://doi.org/10.25934/PR00008428. After pre-processing (see Appendix A2), the final data contain 342 time series from 105 patients. Each time series consists of 54 measurements, taken 5 minutes apart, of a patient’s blood glucose level and exogenous inputs including carbohydrate intake, insulin injection, heart rate and step count from 210 minutes before to 60 minutes after the onset of an exercise instance. We set the first 210 minutes (including 42 time stamps) as historical window and the remaining 60 minutes (including 12 time stamps) as the prediction window.
@@
Evaluation and metrics We adopt repeated CV (see Appendix A4) to estimate various metrics of interest. In addition to all 6 metrics used in synthetic set-ups, we include model variance for measuring robustness and an clinical significance metric, Diagnostic Accuracy–accuracy of classifying a patient as hyper (>180 mg/dl), in-range (80-180 mg/dl) or hypo (<80 mg/dl) using model predictions. All metrics, except Peak RMSE, Peak MAPE and ENP, are reported with 1-sigma SE.
@@
5 Broader impact and limitations
We propose a three-step procedure to simplify the mechanistic graph underlying MNODEs, aiming to improve both predictive accuracy and interpretability. This is especially useful in biomedical domains, where models often involve complex biological processes and high-quality data are limited. By leveraging domain-informed graph refinement, structural pruning, and edge-weight sparsification, our method produces compact and predictive models that align with mechanistic priors while reducing overfitting. This enables more transparent and data-efficient modeling, potentially accelerating discovery in systems biology, personalized medicine, and related fields.
More broadly, our framework contributes to the effort to integrate domain knowledge into deep learning in a principled way. It provides insights into which components (e.g. cycles, delays) of a mechanistic model are predictive, offering guidance for experimental focus and hypothesis generation.
However, key limitations remain. The benefits of our approach diminish in data-rich settings, where black-box models can exploit greater flexibility. It also cannot recover mechanisms omitted from the initial model specification, limiting its use when mechanistic priors are incomplete. Since training is performed on observational data, the resulting simplified models are not causally identifiable and should not inform interventions without further validation.
Finally, while this work focuses on MLP-based decoders, extending the method to other architectures or integrating it into active learning pipelines remains an important direction for future work.
Acknowledgment
This publication is based on research using data from Jaeb Center for Health Research Foundation that has been made available through Vivli, Inc. Vivli has not contributed to or approved, and is not in any way responsible for, the contents of this publication. We thank Alex Wang, Dessi Zeharieva, Emily Fox, Matthew Levine and Ramesh Johari for providing assistance with data acquisition and preliminary discussions on T1D and model reduction.
Appendix

### v2

@@ 1 Introduction
In the field of dynamical system modeling, a significant category of hybrid approaches builds on neural ordinary differential equations (neural ODEs)  (Haber & Ruthotto, 2017; Chen et al., 2018; Kidger, 2021), which arose from the insight that deep residual neural networks can be formulated as continuous-time dynamical systems (Rico-Martinez et al., 1992; Weinan, 2017). Neural ODEs are well-suited for modeling dynamical systems as they offer continuous-time representations with latent dynamics that integrate seamlessly into modern machine learning pipelines with automatic differentiation, making them both scalable and flexible. More recently, researchers have adapted neural ODEs to incorporate domain-knowledge-informed relational inductive bias, often derived from mechanistic models or causal graphs. This hybrid style—sometimes termed Graph Neural ODE (Poli et al., 2019), Neural Causal Model (Xia et al., 2021), Neural State Space Modeling (Hussain et al., 2021) or Mechanistic Neural ODE (MNODE) (Zou et al., 2024)—ensures mechanistic plausibility and interpretability while taking advantage of the flexibility of neural networks, thereby improving model performance and robustness in data-scarce settings.
@@
Traditional model reduction approaches in biochemistry—such as timescale separation (Michaelis & Menten, 1913; Johnson & Goody, 2011) and quasi-steady-state approximations (Schauer & Heinrich, 1983; Bothe & Pierre, 2010)—often require deep domain expertise or extensive trial-and-error. On the other end of the spectrum, data-driven graph-based model reduction offers a pathway to solve this problem. In many healthcare application domains, mechanistic ODEs can be represented as reaction networks or directed graphs, where nodes denote system states and edges denote interactions (Hodgkin & Huxley, 1952; Holz & Fahr, 2001; Smith et al., 2004; Canini & Perelson, 2014; Man et al., 2014). In recent years, many solutions have emerged from the graph neural network (GNN) community for general graph pruning, using approaches such as topology-based node/edge selection (Spielman & Srivastava, 2008; Liu et al., 2023), learning-based sub-graph sampling (Wang et al., 2019; Zeng et al., 2019; Zheng et al., 2020), or optimization-based graph sparsification (Li et al., 2020; Jiang et al., 2021; 2023). However, these reduction methods are typically data-driven and agnostic of any domain knowledge, and thus do not necessarily preserve key mechanistic structures or constraints. Furthermore, non-gradient-based reduction methods (e.g., greedy search) can be prohibitively costly in computation for large, high-dimensional ODE systems. As a result, a gap remains for computationally efficient solutions that reduce model complexity while preserving the mechanistic integrity and improving predictive performance for hybrid neural ODEs.
@@ 2.1 Preliminary
where $S_{\text{pa}(i)}\subseteq S,\,\,X_{\text{pa}(i)}\subseteq X$ are subsets of state and input variables, respectively, on which the derivative of $s_{i}$ with respect to $t$ depends, i.e., they are “parents” of $s_{i}$.
@@
Observable states are variables in the system that can be directly measured through experiments or sensors. These are the quantities that can be collected and tracked over time. On the other hand, latent states are variables that are not directly accessible but still believed to play a role in system dynamics.
@@
2.2 Prediction task: time series forecasting with dynamic covariates
@@
we want to predict $S_{\text{obs}}^{\text{F}}=\{S_{\text{obs}}(t_{k})\}_{k=1}^{q}\in\mathbb{R}^{q\times|S_{\text{obs}}|}$, the future value of the observable states.
@@
2.3 Model architecture: mechanistic neural ODE (MNODE)
@@
$\frac{ds_{i}(t)}{dt}=\text{NN}_{i}(S_{\text{pa}(i)}(t),X_{\text{pa}(i)}(t),t),\quad S(0)=\hat{S}_{\text{lat}}(0)=\text{Encoder}(S^{\text{past}},X^{\text{past}}).$ (1)
@@
2.4 Reduction Algorithm: Hybrid graph sparsification (HGS)
Step 1: merging maximal strongly connected components Given a directed graph representation of a mechanistic ODE system $G=(E,V)$ (since the dependency on the mechanistic ODE system is clear from context, we will omit the $M$ subscript to simplify notations), we first collapse all maximal strongly connected components (MSCCs) in $G$ into super-nodes to make it an RDAG (note that we allow self loops). This is implemented by first partitioning $V$ into disjoint subsets of MSCCs $C_{i}$:
@@ 2.4 Reduction Algorithm: Hybrid graph sparsification (HGS)
Rationale of step 1 Transforming the original graph into an RDAG via collapsing the MSCCs reveals high-level causal structure of the system, simplifies the interpretation, and provides a topological ordering with acyclic structure that improves training stability, as feedback loops are known to cause blow-ups, exploding gradients and stiffness of the ODE system. With cycles, many complicated constraints on the system parameters are needed to simultaneously control stiffness, blow-ups and exploding gradients. Without cycles, the system Jacobian is upper triangular after proper rearrangement, and the corresponding eigenvalues are simply its’ diagonal elements, allowing substantially fewer and simpler constraints on parameters to ensure system stability. (see Appendix A8 for detailed discussions and examples). In addition, replacing each MSCC with a self-loop does not sacrifice much predictive power because neural networks are capable of approximating the effect of complex intra-component dynamics—a key motivation of hybrid modeling Raissi et al. (2019).
Step 1 Customization: While we have chosen the default set-up of HGS step 1 to collapse all MSCCs, users may, based on application needs and their own domain knowledge, choose not to collapse certain MSCCs, and Step 2 and 3 of HGS will still be compatible in these cases. Causal interpretability can be preserved via temporal unfolding, where feedback loops are resolved into time-lagged dependencies between distinct temporal instances of the state variables.
@@
[…]
@@
Intuition and rationale of step 2 Intuitively, one may think of a physiological path as a student’s high-school journey moving through grades 9 to 12. Normally, the student progresses step by step—9 → 10 → 11 → 12. A transitive closure adds all possible “skip-grade” links, letting the student jump directly from grade 9 to 11 or 12, or from 10 to 12, as long as they always move to a higher grade (obey the reachability relations). A partial transitive closure is a more cautious version: it allows some skipping but forbids overly aggressive jumps, like going straight from grade 9 to 12. The idea is that, just as students progress at different speeds, biological processes/pathways in physiological systems also vary in how many intermediate states they pass through and can therefore often be better modeled with fewer latent states. For example, quasi-steady-state approximations in chemical kinetics eliminate fast variables by assuming equilibrium. By adding shortcuts (transitive closure), the model gains flexibility to capture these differences without discarding realistic reachability constraints.
Step 2 Customization Using a partial (rather than full) transitive closure is a choice made by the authors in the context of glucose modeling because it prevents introducing direct input–output edges unsupported by the mechanistic model, and preserving some latent dynamics has been shown to be important (Dalla Man et al., 2009). Similar to step 1, rather than following this default set-up, users may choose to include full transitive closure or omit selected short-cut paths based on their needs.
@@
Step 3 Customization In step 3, while the default set-up is to penalize all edge weights $W$, if certain edges are deemed indispensable or more important by the user, they can be taken out of the regularization term or be given separately a customized $\lambda$.
2.5 Important note: implausibility of true support recovery It should be pointed out that our method is not meant for true support recovery (i.e. recover the true underlying causal graph of the data generating process) because the expressivity of neural networks lead to equivalent MNODE models whose underlying graphs are different (Xia et al., 2021), and the task of recovering the true graph structure is therefore theoretically implausible without making strong assumptions about ground truth. Our goal is to efficiently induce model sparsity and facilitate the generation of data-driven hypotheses, which require further clinical validation, without making strong assumptions about the ground truth other than that it is more sparse than the original mechanistic prior.
@@ 3 Related work
Sparsity, LASSO, and Generalization LASSO-style penalties have long been used to induce sparsity. Applications range from residual networks (Lemhadri et al., 2021), varying-coefficient models (Thompson et al., 2023), and CNNs via group LASSO (Liu et al., 2015; Wen et al., 2016), to feature selection in MLPs (Zhao et al., 2015; Sun et al., 2016; Wang et al., 2017) and local linear sparsity (Ross et al., 2017). However, these methods often yield sparsity patterns that are hard to interpret. Our approach extends LASSO to MNODE while grounding the learned sparsity in mechanistic graph structures, providing interpretability and domain-aligned insights. Prior work has shown that reduced neural networks can generalize as well as, or even better than, their unreduced counterparts (Gale et al., 2019; Bartoldson et al., 2020; Hoefler et al., 2021). For graph-based model reduction, You et al. (2020) found that mildly sparse relational graphs often outperform fully connected ones. Our results support this: MNODEs built from lightly pruned graphs perform better than those using dense graphs in low-data regimes.
@@
Graph Sparsification via Optimization Our method is related to optimization-based graph sparsification approaches for GNNs. For example, Li et al. (2020) constrained the $L_{0}$ norm of the adjacency matrix, Jiang et al. (2021) used elastic net penalties, and Jiang et al. (2023) applied exclusive group LASSO to encourage neighborhood sparsity. Unlike these methods, we do not sparsify the adjacency matrix directly. Instead, we sparsify edge weights in message passing, reducing influence from less informative neighbors. Crucially, GNN sparsification methods are typically data-driven and ignore mechanistic structure. In contrast, our method is structure-aware: it begins with a domain-informed pruning step that restricts the search to physically plausible graphs. This makes our approach more suitable for hybrid models like the Graph Network Simulator (Sanchez-Gonzalez et al., 2020) that require mechanistic consistency. Finally, while Zou et al. (2024) used a greedy, stepwise reduction scheme, our method is more computationally efficient and yields better performance—analogous to the gains of LASSO over stepwise selection in linear models (Hastie et al., 2020).
Landscape of Hybrid Modeling and Trade-offs There are many variants of MNODE in the current landscape of hybrid modeling with different degrees of hybridization and strength of mechanistic prior. In Zou et al. (2024), a more general form of mechanistic neural ODE is proposed as:
$\frac{dS(t)}{dt}=c_{1}f_{m}(S(t),X(t),t;\,\beta(t))+(1-c_{1})f_{nn1}(S(t),X(t),c_{2}Z(t);\,c_{3}A_{m}+(1-c_{3})\mathbf{1})$
$\beta(t)=c_{4}\beta_{m}+(1-c_{4})f_{nn2}(S(t),X(t),t),~~\mbox{and}~~\frac{dZ(t)}{dt}=f_{nn3}(S(t),X(t),Z(t),t),$
where $f_{m},\beta_{m}$ and $A_{m}$ represent the functional form, model parameters and dependency structure (adjacency matrix) of the mechanistic prior, respectively, and $f_{nni},i=1,2$ are neural networks. For $c_{1}=c_{2}=0,\,c_{3}=1$, one recovers MNODE. The deep mechanistic simulator in Miller et al. (2020) uses $c_{1}=1,c_{4}=0$. The neural closure learning model in Gupta & Lermusiaux (2021) uses $c_{1}=0.5,c_{2}=c_{3}=c_{4}=0$. The hybrid ODE model in Qian et al. (2021) uses $c_{1}=0.5,c_{2}=1,c_{3}=0$, while the standard black-box neural ODE uses $c_{1}=c_{3}=0,c_{2}=1$. In general, the more mechanistic components, the more constraints on the function space of the resulting hybrid model. In data-limited regimes, such constraints help the model quickly learn reasonable representations but also undermine the model’s ability to capture complex real-world dynamics not represented by the mechanistic prior. On the other hand, the less mechanistic prior, the more flexible the hybrid model becomes. While the model gains more freedom to learn arbitrary patterns, it is also exposed to risks of overfitting and high variance. The challenge is to find the sweet spot: using just enough mechanistic guidance to keep the model grounded without too much constraining.
@@ 4.1 Experiments on synthetic data
General Setting We assume the data are generated by an unknown sparse dynamical system (ODE), but the mechanistic model is overly complex and contains redundant variables. We show that HGS produces models that are more sparse, predictive and robust than those obtained by existing methods.
@@
Starting graph Assuming the true data generating process is unknown, we consider two settings in which the starting “mechanistic” graph contains redundant structures: (1) a refined graph whose redundant part contains 3 input nodes, 1 latent node, and 1 latent cycle; (2) a comprehensive graph whose redundant part contains 6 input nodes, 3 latent nodes and 3 latent cycles (See Appendix A4.1 for illustrations). All redundant input variables are generated from independent $\mathcal{N}(0,0.5)$.
@@
Evaluation and metrics All evaluations are performed on the test set (see Appendix A5). We report RMSE with 1-sigma standard error (SE) for predictive performance, Peak (worst-case) RMSE for robustness and Effective Number of Parameters (ENP, average number of parameters whose magnitude is $10^{-3}$ under CV-selected hyperparameter setting) for model reduction. Additional metrics including MAPE, Peak MAPE and correlation are reported in A7.1.
@@ 4.2 Experiments on real-world data: blood glucose forecasting for T1D patients
Data Our data are from the T1D Exercise Initiative (T1DEXI)  (Riddell et al., 2023), which is available to public at https://doi.org/10.25934/PR00008428. After pre-processing (see Appendix A3), the final data contain 342 time series from 105 patients. Each time series consists of 54 measurements, taken 5 minutes apart, of a patient’s blood glucose level and exogenous inputs including carbohydrate intake, insulin injection, heart rate and step count from 210 minutes before to 60 minutes after the onset of an exercise instance. We set the first 210 minutes (42 time stamps) as historical window and the remaining 60 minutes (12 time stamps) as the prediction window.
@@
Evaluation and metrics We adopt repeated CV (see Appendix A5) to estimate various metrics of interest. Specifically, we split data randomly at the exercise instance level to evaluate the average predictive performance on both intra- and inter-patient instances, reflecting the likely real-world deployment of the prediction algorithm to both existing and new patients. In addition, since intra-patient variability can be as large as inter-patient variability in T1D modeling (Moscoso-Vasquez et al., 2016; Bell et al., 2021; Laguna et al., 2014), we anticipate similar results from patient-level cross-validation (see Appendix A4.2 for details and empirical evidence). In addition to all 6 metrics used in synthetic set-ups, we include model variance for measuring robustness and an clinical significance metric, Diagnostic Accuracy–accuracy of classifying a patient as hyper ($\geq$ 180 mg/dl), in-range (80-180 mg/dl) or hypo ($\leq$ 80 mg/dl) using model predictions. All metrics, except Peak RMSE, Peak MAPE and ENP, are reported with 1-sigma SE.
@@
Mechanistic interpretation of results HGS can yield interpretable and biologically plausible new insights. For example, HGS chooses to eliminate edges corresponding to glucagon feedback loops, which suggests that impaired glucagon response during hypoglycemia (Seaquist et al., 2013) may also persist during exercise-induced hypoglycemia—a novel hypothesis that could guide future investigations.
@@
5 Broader impact
We propose a three-step procedure to simplify the mechanistic graph underlying MNODEs, aiming to improve prediction, robustness and interpretability. This is especially useful in biomedical domains, where models often involve complex biological processes and high-quality data are limited. By leveraging domain-informed graph refinement, structural pruning, and edge-weight sparsification, our method produces compact and predictive models that align with mechanistic priors while reducing overfitting. This enables more transparent and data-efficient modeling, potentially accelerating discovery in systems biology, personalized medicine, and related fields. More broadly, our framework contributes to the effort to integrate domain knowledge into deep learning in a principled way. It provides insights into which components (e.g. cycles, delays) of a mechanistic model are predictive, offering guidance for experimental focus and hypothesis generation.
Reproducibility Statement
Our data processing pipeline is described in Appendix A3. Model implementation and training details are provided in Appendix A4.2. Model evaluation details are provided in Appendix A5. The mechanistic model used is described in Appendix A6. Our code is also submitted as supplementary materials.
LLM Usage
LLM is only used to aid and polish the writing of the paper. We did not use LLM for any other purpose.
Appendix A Appendix